In a triangle ABC, the measure of angle B is 4 times the measure of angle A and the measure C is 2° less than 5 times the measure of angle A
C=5x-2
B=4x
A=x
Solve for A

Answer by ikleyn(52873)

.

    A + B + C = 180  degrees


    x + 4x + (5x-2) = 180

        10x         = 182

          x         = 182/10 = 18.2 degrees.


ANSWER.  The measure of the angle A is 18.2 degrees.

Solved.

-----------------

In the future, PLEASE, do not post problems with no question.